# Infinitely many hyperelliptic curves of small genus and small fixed rank, and of any genus and rank two

### v3

@@ 2.5. Full rational two-torsion - over any number field
By [Smi22b, Theorem 2.14, Example 3.1], there exists infinitely many (in fact a positive proportion of) square-free elements $m,d\in\mathcal{O}_{K}$ such that $\mathrm{Rank}(E_{1,m}(K))=\mathrm{Rank}(E_{2,d}(K))=0$. Here, we note that there is a subtlety with generating infinitely many such rank 0 quadratic twist families of elliptic curves over any number field $K$. This is where the “parity-invariant subcase” appearing after [Smi22b, Case 2.13] becomes relevant. (We would like to sincerely thank Adam Morgan for pointing out this subtlety.) Our construction of elliptic curves $E_{1,m}$ and $E_{2,d}$ shows that both curves have at least one place of multiplicative reduction. By the first paragraph of proof of [MS, Proposition 4.3] and [MS, Lemma 4.2], performing an unramified quadratic twist at the local condition over such a multiplicative place changes the parity of 2-Selmer group of $E_{1,m}$ (and likewise for $E_{2,d}$). Hence, we can always find infinitely many $m,d\in\mathcal{O}_{K}$ (in fact a positive density of square-free $m,d\in\mathcal{O}_{K}$) such that the parity of 2-Selmer group of $E_{1,m}$ is even. By [Smi22b, Theorem 2.14, Example 3.1], there are infinitely many choices of $m,d\in\mathcal{O}_{K}$ such that $\mathrm{Rank}(\mathrm{Jac}(C_{m,d})(K))=0$.
We turn our focus to constructing infinitely many curves $C_{m,d}$ such that $\mathrm{Rank}(\mathrm{Jac}(C_{m,d})(K))=1$. We note that this result for $g=2$ was obtained before [KM]. By our choice of nine prime elements $p_{i},q_{j},r_{k}$ the elliptic curve $E_{1,1}$ satisfies the following three conditions:
@@ 3. Higher genus curves
We generalise the method of gluing two elliptic curves explored in Section 2 to gluing two hyperelliptic curves. This allows us to construct infinitely many hyperelliptic curves of any genus over any number field whose Jacobians have fixed positive rank between 0 and $2$. We also explore examples of genus 3, 4, 5, and 6 hyperelliptic curves over $\mathbb{Q}$ with fixed small positive rank.
3.1. Higher genus - over any number field
Let $K$ be a number field, $f=x^{n}+a_{n-1}x^{n-1}+\cdots+a_{1}x+a_{0}\in\mathcal{O}_{K}[x]$ a monic polynomial of degree $n$, and $d\in\mathcal{O}_{K}$. Define
@@ 3.1. Higher genus - over any number field
i.e., the polynomial such that $f_{d}(dx)=d^{n}f(x)$ Consider the following three hyperelliptic curves
$C$ $\colon dy^{2}=f_{d}(x^{2});$
$C_{1}$ $\colon dy^{2}=xf_{d}(x);$
$C_{2}$ $\colon dy^{2}=f_{d}(x).$
@@
Lemma 3.1.
The map $\Psi=({\psi_{1}}_{*},{\psi_{2}}_{*})\colon\mathrm{Jac}(C)\longrightarrow\mathrm{Jac}(C_{1})\times\mathrm{Jac}(C_{2})$ is an isogeny defined over $K$.
Proof.
The map $\Psi$ is defined over $K$ as $\psi_{1},\psi_{2}$ are defined over $K$. We denote by $C^{(g(C))}$, $C_{1}^{(g(C_{1}))}$, and $C_{2}^{(g(C_{2}))}$ the symmetric powers of $C$, $C_{1}$, and $C_{2}$ by respective genus g(C)=n1g(C)=n-1, $g(C_{1})=\lfloor\frac{n}{2}\rfloor$ and $g(C_{2})=\lfloor\frac{n-1}{2}\rfloor$. Using [Mil86b, Theorem 5.1, Corollary 7.5], we have the following commutative diagram.
$C^{g}$$C_{1}^{g(C_{1})}\times C_{2}^{g(C_{2})}$$C^{(g(C))}$$C_{1}^{(g(C_{1}))}\times C_{2}^{(g(C_{2}))}$$\mathrm{Jac}(C)$$\mathrm{Jac}(C_{1})\times\mathrm{Jac}(C_{2})$$\scriptstyle{\psi_{1}\times\psi_{2}}$$\scriptstyle{\Psi}$
The surjectivity of $\Psi$ follows from the surjectivity of $C^{g}\rightarrow C_{1}^{g(C_{1})}\times C_{2}^{g(C_{2})}$ and all vertical maps in the commutative diagram. Because $\Psi$ is a homomorphism of abelian varieties and since $\lfloor\frac{n}{2}\rfloor+\lfloor\frac{n-1}{2}\rfloor=n-1$, we have $\dim\mathrm{Jac}(C)=\dim(\mathrm{Jac}(C_{1})\times\mathrm{Jac}(C_{2}))$, so, we can use [Mil86a, Proposition 8.1] to show that $\Psi$ is an isogeny. ∎
Remark 3.2.
Thanks to Maarten Derickx’s suggestion, we can note that the pull-backs $\psi_{1}^{*}(\mathrm{Jac}(C_{1}))\subseteq\mathrm{Jac}(C)$ and $\psi_{2}^{*}(\mathrm{Jac}(C_{2}))\subseteq\mathrm{Jac}(C)$ have 0-dimensional intersection inside $\mathrm{Jac}(C)$. This can be checked by pulling-back the holomorphic differentials $\frac{dx}{y},\ldots\frac{x^{\lfloor\frac{n}{2}\rfloor-1}dx}{y}$ from $\mathrm{Jac}(C_{1})$ and $\frac{dx}{y},\ldots\frac{x^{\lfloor\frac{n-1}{2}\rfloor-1}dx}{y}$ from $\mathrm{Jac}(C_{2})$, which give the full basis of the holomorphic differentials $\frac{dx}{y},\ldots\frac{x^{n-2}dx}{y}$ on $\mathrm{Jac}(C)$.
Note that, depending on the parity of $n$, after a linear change of variables ($x\mapsto dx$), we can simplify the equations for $C_{1}$ and $C_{2}$
$C_{1}\cong y^{2}=xf(x),\;C_{2}\cong dy^{2}=f(x)$ $\text{ if }2\mid n,$
$C_{1}\cong dy^{2}=xf(x),\;C_{2}\cong y^{2}=f(x)$ $\text{ if }2\nmid n.$
Note that in both cases, one of the curves is isomorphic to a fixed one, given by an odd degree model, whereas an even degree model becomes a twist by $d$. Hence, this gives a clear strategy to construct infinitely many curves of genus $g$ of a fixed rank $r$; we just need to find a specific one satisfying some mild conditions.
@@
  • For every $1\leq i\leq 2g$, the place $\omega^{\prime}_{i}$ is of type $\{i,2g+1\}$.
@@
We focus on understanding the Mordell-Weil rank of the family of quadratic twists of hyperelliptic curves that are BB-über-generic. Before doing so, we recall the following condition on $C$ from [BM, Condition 5.1] and [MS, Proposition 4.3]. Once again, we would like to thank Adam Morgan for suggesting the importance of the third condition provided below.
Condition 3.5 (Condition 5.1 from [BM] and Proposition 4.3 from [MS]).
@@
We note that if $C$ is $B$-generic, then there exists a square-free element $d\in\mathcal{O}_{K}^{\times}$ such that $C_{d}$ satisfies Condition 3.5. To see this, we note that [BM, Proposition 8.10] shows that any hyperelliptic curve $C:y^{2}=(x-a_{1})\cdots(x-a_{2g+1})$ satisfying the following two conditions
@@
satisfies the first two requirements in Condition 3.5. By [KM, Theorem 3.8], there exists a square-free element $d\in\mathcal{O}_{K}^{\times}$ such that the quadratic twist $C_{d}$ of a $B$-generic hyperelliptic curve $C$ satisfies the two bullet points mentioned above. By the condition on $a_{i}$’s, one can easily check that $C$ has a single nodal singularity at the place $\omega_{1}$, which satisfies the last condition from Condition 3.5.
@@
Let $C$ be a hyperelliptic curve over $\mathbb{Q}$. One can check the first two conditions from Condition 3.5 by computing the Weil pairing $e_{2}:\mathrm{Jac}(C)[2]\times\mathrm{Jac}(C)[2]\to\mu_{2}$. Suppose $\langle P_{1},P_{2},\cdots,P_{2g}\rangle$ is a basis for the set $\mathrm{Jac}(C)[2]$. If for every pair $1\leq i<j\leq 2g$ we have $e_{2}(P_{i},P_{j})=-1$, then the first two conditions hold by [BM, Proposition 2.15].
For the special case when $C$ is given by the Weierstrass model $C:y^{2}=(x-b_{1})\cdots(x-b_{4k+3})$ for some positive integer $k$, we can combine [Yu16, Corollary 7.6] and [Smi22b, Theorem 2.14] to ensure that at least $50\%$ of Jacobians of quadratic twists of $C$ have rank 0. In the case when $C$ is not of the form above, we can replace the third condition in Condition 3.5 with the requirement that there exists a single square-free element $d\in\mathbb{Q}^{\times}$ such that $\dim_{\mathbb{F}_{2}}\mathrm{Sel}_{2}(\mathrm{Jac}(C_{d})/\mathbb{Q})\equiv 0\mod 2$. This condition ensures that there is a positive density of Jacobians of rank 0, as shown in the discussion following after [Smi22b, Case 2.13] and the main theorem of [BM].
@@
The proof of [KM, Corollary 6.15] and the first paragraph of the proof of [MS, Proposition 4.3] implies that all the conditions in [Smi22b, Case 2.13] are satisfied, and that a positive density of square-free elements $d\in\mathcal{O}_{K}$ satisfies $\dim_{\mathbb{F}_{2}}\mathrm{Sel}_{2}(\mathrm{Jac}(C_{d})/K)\equiv 0\mod 2$. Hence, we can apply [Smi22b, Theorem 2.14] to show that there are infinitely many quadratic twists of $C$ such that $\mathrm{Rank}(\mathrm{Jac}(C_{d})/K)=0$. ∎
@@
of rank $r$ as $\mathrm{Jac}(C_{d})$ is isogenous to $\mathrm{Jac}(X)\times\mathrm{Jac}(Y_{d})$ by Lemma 3.1.
@@
of rank 0. By Lemma 3.1, there are infinitely many curves
@@
    Assume that g=2k1g=2k-1 for some $k\in\mathbb{N}$ and that there exists a hyperelliptic curve $X:y^{2}=x(x-1)f(x)$ such that $\mathrm{Rank}(\mathrm{Jac}(X)(K))=r$, where $f\in\mathcal{O}_{K}[x]$, $\deg(f)=2k-1$, is a monic polynomial whose splitting field has Galois group isomorphic to $S_{2k-1}$. Suppose in addition that there exists a square-free element $m\in\mathcal{O}_{K}$ such that the curve $Y_{m}:my^{2}=(x-1)f(x)$ satisfies $\dim_{\mathbb{F}_{2}}\mathrm{Sel}_{2}(\mathrm{Jac}(Y_{m})/K)\equiv 0\mod 2$. Then, there are infinitely many many genus $g$ curves $C_{d}:dy^{2}=(x^{2}-d^{2})f_{d}(x^{2})$ of rank $r$, where $d\in\mathcal{O}_{K}$.
@@
    We note that the existence of a square-free element $m\in\mathcal{O}_{K}$ such that $Y_{m}:my^{2}=xf(x)$ satisfies $\dim_{\mathbb{F}_{2}}\mathrm{Sel}_{2}(\mathrm{Jac}(Y_{m})/K)\equiv 0\mod 2$ implies that the disparity $\delta(\mathrm{Jac}(Y_{1})/K)\neq-\frac{1}{2}$ appearing in the statement of [Yu19, Theorem 1.2]. The assumptions assure that we can apply [Yu19, Theorem 1.2] to ensure that there are infinitely many quadratic twists $Y_{d}\colon dy^{2}=xf(x)$ of $Y\colon y^{2}=xf(x)$ such that $\mathrm{Rank}(\mathrm{Jac}(Y_{d})(K))=0$. Then, the statement follows from Lemma 3.1 which gives that $\mathrm{Jac}(C_{d})$ is isogenous to $\mathrm{Jac}(X)\times\mathrm{Jac}(Y_{d})$ over $K$, and the argument that there are infinitely many curves that are non-isomorphic among them is as before.
@@
    The same as (1), and now we apply [Yu19, Theorem 1.2] to conclude that there are infinitely many quadratic twists $Y_{d}\colon dy^{2}=(x-1)f(x)$ of $Y\colon y^{2}=(x-1)f(x)$ such that $\mathrm{Rank}(\mathrm{Jac}(Y_{d})(K))=0$.
@@
Let $k$ be a positive integer. There is $B>0$ such that the following holds. Let $X:y^{2}=(x-a_{1})\cdots(x-a_{2k+1})$ be a hyperelliptic curve over a number field $K$ that is BB-über-generic such that $\mathrm{Rank}(\mathrm{Jac}(C)(K))=r$. Then there are infinitely many hyperelliptic curves $X/K$ of genus $g$, where $g\in\{2k-1,2k\}$ satisfying $\mathrm{Rank}(\mathrm{Jac}(C)(X))=r+1$.

### v4

@@ 2.5. Full rational two-torsion - over any number field
By [SMI22b, Theorem 2.14, Example 3.1], there exists infinitely many (in fact a positive proportion of) square-free elements $m,d\in\mathcal{O}_{K}$ such that $\mathrm{Rank}(E_{1,m}(K))=\mathrm{Rank}(E_{2,d}(K))=0$. Here, we note that there is a subtlety with generating infinitely many such rank 0 quadratic twist families of elliptic curves over any number field $K$. This is where the “parity-invariant subcase” appearing after [SMI22b, Case 2.13] becomes relevant. (We would like to sincerely thank Adam Morgan for pointing out this subtlety.) Our construction of elliptic curves $E_{1,m}$ and $E_{2,d}$ shows that both curves have at least one place of multiplicative reduction. By the first paragraph of proof of [31, Proposition 4.3] and [31, Lemma 4.2], performing an unramified quadratic twist at the local condition over such a multiplicative place changes the parity of 2-Selmer group of $E_{1,m}$ (and likewise for $E_{2,d}$). Hence, we can always find infinitely many $m,d\in\mathcal{O}_{K}$ (in fact a positive density of square-free $m,d\in\mathcal{O}_{K}$) such that the parity of 2-Selmer group of $E_{1,m}$ is even. By [SMI22b, Theorem 2.14, Example 3.1], there are infinitely many choices of $m,d\in\mathcal{O}_{K}$ such that $\mathrm{Rank}(\mathrm{Jac}(C_{m,d})(K))=0$.
We turn our focus to constructing infinitely many curves $C_{m,d}$ such that $\mathrm{Rank}(\mathrm{Jac}(C_{m,d})(K))=1$. We note that this result for $g=2$ was obtained before [23]. By our choice of nine prime elements $p_{i},q_{j},r_{k}$ the elliptic curve $E_{1,1}$ satisfies the following three conditions:
@@ 3. Higher genus curves
We now use a more general construction than the method of glueing two elliptic curves explored in Section 2, namely, we produce hyperelliptic curves, which are the fibre product of two hyperelliptic curves. This allows us to construct infinitely many hyperelliptic curves of any genus over any number field whose Jacobians have fixed positive rank between 0 and $2$. We also explore examples of genus 3, 4, 5, and 6 hyperelliptic curves over $\mathbb{Q}$ with fixed small positive rank.
3.1. Fibre product of hyperelliptic curves
Let $K$ be a number field, $f,g\in K[x]$ different squarefree polynomials, and consider hyperelliptic curves $C_{f}\colon y^{2}=f(x)$ and $C_{g}\colon y^{2}=g(x)$. Then, there are double covers $\pi_{f}\colon C_{f}\longrightarrow\mathbb{P}^{1}$ and $\pi_{g}\colon C_{g}\longrightarrow\mathbb{P}^{1}$. Hence, there is a curve $C/K$ whose function field $K(C)$ is the compositum $K(C_{f})\otimes_{K(x)}K(C_{g})$. If we denote $L=K(x)$ (the function field of $\mathbb{P}^{1}$), then $K(C_{f})=L(\sqrt{f(x)})$ and $K(C_{g})=L(\sqrt{g(x)})$, and it follows that $K(C)=L(\sqrt{f(x)},\sqrt{g(x)})$. Then, $C$ is the fibre product of $C_{f}$ and $C_{g}$ over $\mathbb{P}^{1}$. Furthermore, we note that $K(C)=L(\sqrt{f(x)},\sqrt{g(x)})=L(\sqrt{f(x)},\sqrt{f(x)g(x)})=L(\sqrt{f(x)g(x)},\sqrt{g(x)})$. So, if we denote $h(x)$ to be the squarefree part of $f(x)g(x)$, then $C$ can also be viewed as the fibre product of $C_{f}$ and $C_{h}\colon y^{2}=h(x)$ over $\mathbb{P}^{1}$, and similarly also as the fibre product of $C_{g}$ and $C_{h}$.
Using the functoriality of Jacobians, the maps $\pi_{f},\pi_{g},\pi_{h}:C\longrightarrow C_{f},C_{g},C_{h}$ induce surjective homomorphisms $\psi_{f}:\mathrm{Jac}(C)\longrightarrow\mathrm{Jac}(C_{f})$, $\psi_{g}:\mathrm{Jac}(C)\longrightarrow\mathrm{Jac}(C_{g})$, and $\psi_{h}:\mathrm{Jac}(C)\longrightarrow\mathrm{Jac}(C_{h})$, see [MUM74, p.328] for a proof for surjectivity of these maps. We obtain a homomorphism $\psi=(\psi_{f},\psi_{g},\psi_{h})\colon\mathrm{Jac}(C)\longrightarrow\mathrm{Jac}(C_{f})\times\mathrm{Jac}(C_{g})\times\mathrm{Jac}(C_{h})$ defined over $K$. We will show that it is surjective as well. Recall that the Prym variety $\mathrm{Prym}(C/C_{f})$ is the kernel of $\psi_{f}$, because the function field $K(C)/K(C_{f})$ is ramified at places dividing $h(x)$. and we have $\mathrm{Jac}(C)\sim\mathrm{Jac}(C_{f})\times\mathrm{Prym}(C/C_{f})$, which is defined over $K$. By [MUM74, page 346, statement (c)], we have an isomorphism $\mathrm{Prym}(C/C_{f})\cong\mathrm{Jac}(C_{g})\times\mathrm{Jac}(C_{h})$. Hence, one obtains an isogeny between $\mathrm{Jac}(C)$ and $\mathrm{Jac}(C_{f})\times\mathrm{Jac}(C_{g})\times\mathrm{Jac}(C_{h})$ over $\overline{K}$; in particular, this map is surjective over $\overline{K}$. This implies that $\dim(\mathrm{Jac}(C))=\dim(\mathrm{Jac}(C_{f}))+\dim(\mathrm{Jac}(C_{g}))+\dim(\mathrm{Jac}(C_{h}))$, so $g(C)=g(C_{f})+g(C_{g})+g(C_{h})$, where $g(X)$ denotes the genus of the curve $X$. We use a slight abuse of notation with $g$ being a genus and a polynomial, but we believe that it is clear in the context which one is which. Given a smooth projective curve $X$, we denote by $X^{k}$ the product of $k$ copies of $X$, and by $X^{(k)}$ the $k$-th symmetric power of $X$. By using [MIL86b, Theorem 5.1, Corollary 7.5], we have the following commutative diagram.
$C^{g(C)}$$C_{f}^{g(C_{f})}\times C_{g}^{g(C_{g})}\times C_{h}^{g(C_{h})}$$C^{(g(C))}$$C_{f}^{(g(C_{f}))}\times C_{g}^{(g(C_{g}))}\times C_{h}^{(g(C_{h}))}$$\mathrm{Jac}(C)$$\mathrm{Jac}(C_{f})\times\mathrm{Jac}(C_{g})\times\mathrm{Jac}(C_{h})$$\scriptstyle{\pi_{f}^{g(C_{f})}\times\pi_{g}^{g(C_{g})}\times\pi_{h}^{g(C_{h})}}$$\scriptstyle{\pi_{f}^{g(C_{f})}\times\pi_{g}^{g(C_{g})}\times\pi_{h}^{g(C_{h})}}$$\scriptstyle{\psi}$
The upper right arrow map is clearly surjective as $g(C)=g(C_{f})+g(C_{g})+g(C_{h})$, and all vertical maps are also surjective, hence, the bottom right map ($\psi$) is also surjective (over $\overline{K}$). Now, [MIL86a, Proposition 7.1] implies that $\psi$ is an isogeny. Since $\psi$ is defined over $K$, we have proven the following lemma.
Lemma 3.1.
The map $\psi\colon\mathrm{Jac}(C)\longrightarrow\mathrm{Jac}(C_{f})\times\mathrm{Jac}(C_{g})\times\mathrm{Jac}(C_{h})$ is an isogeny defined over $K$.
3.2. Higher genus - our strategy
Lemma 3.1 gives an idea on how to look for curves of any genus $g$ of some prescribed rank by descending to the same question but for smaller ranks of curves of smaller genus.
Given our strategy in Section 2, we can try to consider twists of hyperelliptic curves $C_{f}$, $C_{g}$, and $C_{h}$. However, their equations are related. If we choose to twist $C_{f}$ by $d$, and $C_{g}$ to twist by $m$ (we also allow $m=1$, i.e., preserve $C_{g}$), then in our construction, we need to consider the twist of $C_{h}$ by $dm$. Therefore, even though we can make quadratic twists of $C_{f}$ and $C_{g}$ to be independent, the quadratic twist on $C_{h}$ is not anymore independent. It will happen only in the case when $m=d$, but then the twists of $C_{f}$ and $C_{g}$ are not independent. Hence, to study the $\mathrm{Rank}(J(C)(K))$ as the sum of the ranks of Jacobians the twists of $C_{f}$, $C_{g}$, and $C_{h}$ is an extremely difficult problem in general because we need to control simultaneously dependent twists of different curves.
Thus, our idea, which makes this strategy more efficient, is to make sure that one of the curves has genus 0, and then we can simply ignore its Jacobian. We achieve it by having a concrete relation between polynomials $f$ and $g$; namely, we require $f(x)=\gamma xg(x)$, where $\gamma\in K$ is a non-zero constant, and $g(0)\neq 0$ so that we consider smooth curves. Then, $C_{h}\colon y^{2}=\gamma x$ is isomorphic to $\mathbb{P}^{1}$ and will not contribute to the rank of the Jacobian.
Furthermore, in this case, we obtain that the fibre product curve $C$ is a hyperelliptic curve and we get its explicit equation. Looking at the compositum of the function fields of $C_{f}$ and $C_{g}$, we have $K(C)=L(\sqrt{f(x)},\sqrt{g(x)})=L(\sqrt{x},\sqrt{g(x)})$. Now, we can identify the field $L(\sqrt{x})$ with $K(t)$ by putting $t=\sqrt{x}$, so it can also be seen as the function field of $\mathbb{P}^{1}$. Under this identification, we have that $L(\sqrt{x},\sqrt{g(x)})\cong K(t,\sqrt{g(t^{2})})$. The curve $C:y^{2}=g(x^{2})$ has the function field $K(t,\sqrt{g(t^{2})})$, hence $C$ is isomorphic to the fibre product of $C_{f}$ and $C_{g}$, and we may identify them.
[…]
@@ 3.2. Higher genus - our strategy
i.e., the polynomial such that $f_{d}(dx)=d^{n}f(x)$.
Remark 3.2.
The fibre product construction in this case can be seen also very explicitly using the equations of the curves and the maps between them. Consider the following three hyperelliptic curves
$C$ $\colon my^{2}=f_{d}(x^{2});$
$C_{1}$ $\colon my^{2}=xf_{d}(x);$
$C_{2}$ $\colon my^{2}=f_{d}(x).$
@@
We know (by what we proved in §3.1) that $C$ is the fibre product of $C_{1}$ and $C_{2}$ over $\mathbb{P}^{1}$ and $\psi=({\psi_{1}}_{*},{\psi_{2}}_{*})\colon\mathrm{Jac}(C)\longrightarrow\mathrm{Jac}(C_{1})\times\mathrm{Jac}(C_{2})$ is an isogeny defined over $K$.
One can note that $C$ has a group of automorphisms that contains $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ as the maps $(x,y)\mapsto(\pm x,\pm y)$ are automorphisms of $C$. And the three quotients of $C$ by these nontrivial automorphisms are precisely $C_{1}$, $C_{2}$, and $\mathbb{P}^{1}$. Hence, we could use the work of Kani and Rosen [KR89] to conclude that $\mathrm{Jac}(C)\sim\mathrm{Jac}(C_{1})\times\mathrm{Jac}(C_{2})$.
Additionally, there is another way to check that $\psi$ is an isogeny, as suggested by Maarten Derickx. We can note that the pull-backs $\psi_{1}^{*}(\mathrm{Jac}(C_{1}))\subseteq\mathrm{Jac}(C)$ and $\psi_{2}^{*}(\mathrm{Jac}(C_{2}))\subseteq\mathrm{Jac}(C)$ have 0-dimensional intersection inside $\mathrm{Jac}(C)$. This can be checked by pulling-back the holomorphic differentials $\frac{dx}{y},\ldots\frac{x^{\lfloor\frac{n}{2}\rfloor-1}dx}{y}$ from $\mathrm{Jac}(C_{1})$ and $\frac{dx}{y},\ldots\frac{x^{\lfloor\frac{n-1}{2}\rfloor-1}dx}{y}$ from $\mathrm{Jac}(C_{2})$, which give the full basis of the holomorphic differentials $\frac{dx}{y},\ldots\frac{x^{n-2}dx}{y}$ on $\mathrm{Jac}(C)$.
Finally, we can explain what is our strategy to construct hyperelliptic curves of any genus and fixed rank between 0 and 2 (or conditional ranks under the existence of a curve of a specific shape and rank). Note that, depending on the parity of $n$, after a linear change of variables ($x\mapsto dx$), we can simplify the equations for $C_{1}$ and $C_{2}$
(4) $C_{1}\cong dmy^{2}=xf(x),\;C_{2}\cong my^{2}=f(x)$ $\text{ if }2\mid n,$
(5) $C_{1}\cong my^{2}=xf(x),\;C_{2}\cong dmy^{2}=f(x)$ $\text{ if }2\nmid n.$
In general, we consider two twists of curves $C_{1}$ and $C_{2}$, but after specifying $m$, the other twist is independent of $m$ (by choosing an appropriate $d$). This is essential for the proof of Theorem 3.13.
However, for the theorems before, we can just consider the case when $m=d$, when, one of the curves is isomorphic to a fixed one, given by an odd degree model, whereas an even degree model becomes a twist by $d$. In this case, this gives a clear strategy to construct infinitely many curves of genus $g$ of a fixed rank $r$; we just need to find a specific one satisfying some mild conditions.
3.3. Our constructions
@@
  • For every $1\leq i\leq 2g$, the place $\omega^{\prime}_{i}$ is multiplicative of type $\{i,2g+1\}$.
@@
We focus on understanding the Mordell-Weil rank of the family of quadratic twists of hyperelliptic curves that are BB-über-generic. Before doing so, we recall the following condition on $C$ from [4, Condition 5.1] and [31, Proposition 4.3]. Once again, we would like to thank Adam Morgan for suggesting the importance of the third condition provided below.
Condition 3.5 (Condition 5.1 from [4] and Proposition 4.3 from [31]).
@@
We note that if $C$ is $B$-generic, then there exists a square-free element $d\in\mathcal{O}_{K}^{\times}$ such that $C_{d}$ satisfies Condition 3.5. To see this, we note that [4, Proposition 8.10] shows that any hyperelliptic curve $C:y^{2}=(x-a_{1})\cdots(x-a_{2g+1})$ satisfying the following two conditions
@@
satisfies the first two requirements in Condition 3.5. By [23, Theorem 3.8], there exists a square-free element $d\in\mathcal{O}_{K}^{\times}$ such that the quadratic twist $C_{d}$ of a $B$-generic hyperelliptic curve $C$ satisfies the two bullet points mentioned above. By the condition on $a_{i}$’s, one can easily check that $C$ has a single nodal singularity at the place $\omega_{1}$, which satisfies the last condition from Condition 3.5.
@@
Let $C$ be a hyperelliptic curve over $\mathbb{Q}$. One can check the first two conditions from Condition 3.5 by computing the Weil pairing $e_{2}:\mathrm{Jac}(C)[2]\times\mathrm{Jac}(C)[2]\to\mu_{2}$. Suppose $\langle P_{1},P_{2},\cdots,P_{2g}\rangle$ is a basis for the set $\mathrm{Jac}(C)[2]$. If for every pair $1\leq i<j\leq 2g$ we have $e_{2}(P_{i},P_{j})=-1$, then the first two conditions hold by [4, Proposition 2.15].
For the special case when $C$ is given by the Weierstrass model $C:y^{2}=(x-b_{1})\cdots(x-b_{4k+3})$ for some positive integer $k$, we can combine [YU16, Corollary 7.6] and [SMI22b, Theorem 2.14] to ensure that at least $50\%$ of Jacobians of quadratic twists of $C$ have rank 0. In the case when $C$ is not of the form above, we can replace the third condition in Condition 3.5 with the requirement that there exists a single square-free element $d\in\mathbb{Q}^{\times}$ such that $\dim_{\mathbb{F}_{2}}\mathrm{Sel}_{2}(\mathrm{Jac}(C_{d})/\mathbb{Q})\equiv 0\mod 2$. This condition ensures that there is a positive density of Jacobians of rank 0, as shown in the discussion following after [SMI22b, Case 2.13] and the main theorem of [4].
@@
The proof of [23, Corollary 6.15] and the first paragraph of the proof of [31, Proposition 4.3] implies that all the conditions in [SMI22b, Case 2.13] are satisfied, and that a positive density of square-free elements $d\in\mathcal{O}_{K}$ satisfies $\dim_{\mathbb{F}_{2}}\mathrm{Sel}_{2}(\mathrm{Jac}(C_{d})/K)\equiv 0\mod 2$. Hence, we can apply [SMI22b, Theorem 2.14] to show that there are infinitely many quadratic twists of $C$ such that $\mathrm{Rank}(\mathrm{Jac}(C_{d})/K)=0$. ∎
@@
of rank $r$ as $\mathrm{Jac}(C_{d})$ is isogenous to $\mathrm{Jac}(X)\times\mathrm{Jac}(Y_{d})$ by Remark 3.2.
@@
of rank 0. By Remark 3.2, there are infinitely many curves
@@
    Assume that g=2k1g=2k-1 for some $k\in\mathbb{N}$ and that there exists a hyperelliptic curve $X:y^{2}=x(x-1)f(x)$ such that $\mathrm{Rank}(\mathrm{Jac}(X)(K))=r$, where $f\in\mathcal{O}_{K}[x]$, $\deg(f)=2k-1$, is a monic polynomial whose splitting field has Galois group isomorphic to $S_{2k-1}$. Suppose in addition that there exists a square-free element $m\in\mathcal{O}_{K}$ such that the curve $Y_{m}:my^{2}=(x-1)f(x)$ satisfies $\dim_{\mathbb{F}_{2}}\mathrm{Sel}_{2}(\mathrm{Jac}(Y_{m})/K)\equiv 0\mod 2$. Then, there are infinitely many genus $g$ curves $C_{d}:dy^{2}=(x^{2}-d^{2})f_{d}(x^{2})$ of rank $r$, where $d\in\mathcal{O}_{K}$.
@@
    We note that the existence of a square-free element $m\in\mathcal{O}_{K}$ such that $Y_{m}:my^{2}=xf(x)$ satisfies $\dim_{\mathbb{F}_{2}}\mathrm{Sel}_{2}(\mathrm{Jac}(Y_{m})/K)\equiv 0\mod 2$ implies that the disparity $\delta(\mathrm{Jac}(Y_{1})/K)\neq-\frac{1}{2}$ appearing in the statement of [YU19, Theorem 1.2]. The assumptions assure that we can apply [YU19, Theorem 1.2] to ensure that there are infinitely many quadratic twists $Y_{d}\colon dy^{2}=xf(x)$ of $Y\colon y^{2}=xf(x)$ such that $\mathrm{Rank}(\mathrm{Jac}(Y_{d})(K))=0$. Then, the statement follows from Remark 3.2 which gives that $\mathrm{Jac}(C_{d})$ is isogenous to $\mathrm{Jac}(X)\times\mathrm{Jac}(Y_{d})$ over $K$, and the argument that there are infinitely many curves that are non-isomorphic among them is as before.
@@
    The proof is the same as in (1), and now we apply [YU19, Theorem 1.2] to conclude that there are infinitely many quadratic twists $Y_{d}\colon dy^{2}=(x-1)f(x)$ of $Y\colon y^{2}=(x-1)f(x)$ such that $\mathrm{Rank}(\mathrm{Jac}(Y_{d})(K))=0$.
@@
Let $k>1$ be a positive integer. There is $B>0$ such that the following holds. Let $X:y^{2}=(x-a_{1})\cdots(x-a_{2k+1})$ be a hyperelliptic curve over a number field $K$ that $X$ is BB-über-generic and it has a twist $X_{m}:my^{2}=(x-a_{1})\cdots(x-a_{2k+1})$ such that $\mathrm{Rank}(\mathrm{Jac}(X_{m})(K))=r$. Then there are infinitely many hyperelliptic curves $C/K$ of genus $g$, where $g\in\{2k-1,2k\}$ satisfying $\mathrm{Rank}(\mathrm{Jac}(C)(K))=r+1$.